40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 25 10:48:13 2008, Koji, Update, General, Abs. Len. Meas. ~ Realignment / beating / PLL trial 
    Reply  Mon Jul 28 15:54:04 2008, Koji, Update, General, Abs. Len. Meas. ~ More on the beat / the PLL setup PLL_setup.pngbeat_at_8MHz.jpg
       Reply  Thu Jul 31 11:19:30 2008, Koji, Update, General, Abs. Len. Meas. ~ Resonance search trial 
          Reply  Sat Aug 2 12:53:43 2008, Koji, Update, General, Abs. Len. Meas. ~ New PLL at the PSL table yarm_dc.png
             Reply  Mon Aug 4 00:37:58 2008, Koji, Update, General, Abs. Len. Meas. ~ RF PD at the Y end / Manual frequency scan YEND_LAYOUT.png4th_FSR1.png4th_FSR2.png
                Reply  Mon Aug 4 05:23:57 2008, Koji, Update, General, Abs. Len. Meas. ~ Measurement for Y-arm completed 
                Reply  Mon Aug 4 21:48:24 2008, Koji, Update, General, Abs. Len. Meas. ~ Scan for TEM00/01/10 TEM00.pngTEM01.pngknife_edge.png
                   Reply  Mon Aug 4 22:31:10 2008, Koji, Update, General, Abs. Len. Meas. ~ Simple Test for TEM01/10 split TEM01_10.png
                   Reply  Tue Aug 5 00:05:57 2008, Koji, Update, General, Abs. Len. Meas. ~ IFR2023A calibration 
                      Reply  Wed Aug 6 11:10:34 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM00 scan TEM00fit.pngTEM00FSRfit.png
                         Reply  Wed Aug 6 11:43:52 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM01 scan TEM01fit.png
                            Reply  Thu Aug 14 21:52:51 2008, Koji, Summary, General, Abs. Len. Meas. ~ summary of my Summer mode_spacing_measurement_080816_v2.pdf
Message ID: 787     Entry time: Mon Aug 4 00:37:58 2008     In reply to: 782     Reply to this: 789   793
Author: Koji 
Type: Update 
Category: General 
Subject: Abs. Len. Meas. ~ RF PD at the Y end / Manual frequency scan 
Work log on August 2nd

o Just remind you:
The idea of the absolute length measurement was to detect an RF beat between the injection beam and the PSL beam by resonating both of the beams to the cavity at the same time, but on different londitudinal modes. From the frequency separation between the two beams, we get the FSR of the cavity. In order to have an injection beam with stable frequency separation, a heterodne interferometer was built at the PSL table, and the PLL servo is used to control and stabilize the frequency of the inj. beam.

----------

o An RF PD (Tholab PDA255) and a steering mirror were placed at the Yarm END. Fortunately, I found that an unused BS was already in the optical path. There was a beam block which dump the reflection of the BS and some stray lights of the OPLEV. I moved the beam block to make the BS reflection available, as well as to block the OPLEV stray light still (Photo1). In order to have the RF signal from the PD, a long BNC cable was laid along the Yarm. I did't know any better idea than this. Don't blame me.

o To have an intuitive interpretation of the beat frequency, the injection beam was set to be at higher frequency than the PSL beam. How did I confirm this? When the crystal temp (LT) of the NPRO was tuned to be higher, the beat frequency got lower.

o Frequency of the PLL was manually swept at around 15.51MHz where the 4th FSR was expected to be found. I could see strong RF peak at that frequency! When I tuned the PLL frequency, the peak height changes dramatically! Too cool!

o The amplitude of the RF peak was measured by an RF spectrum analyzer. I did all of this scan by my hands and eyes. The center frequency of the 4th FSR was 15.5149MHz. From the eye I would say the error is +/-150Hz. It is OK so far although I am not sure statistically this is correct or not. This corresponds to the length of 38.64575 +/- 0.00037 [m].

o All of the past measurements are fairly consistent.
Y-arm length
e-log              length [m]           Measurement Conditions
----------------------------------------------------------------------------------------
 556(2008-Jun-24)  38.67    +/- 0.03    Cavity swinging measurement
 776(2008-Jul-31)  38.640   +/- 0.007   Beam injection, poor PLL, Transmitted DC
 782(2008-Aug-02)  38.6455  +/- 0.0012  Beam injection, independent PLL, Transmitted DC
this(2008-Aug-04)  38.64575 +/- 0.00037 Beam injection, independent PLL, Transmitted RF
----------------------------------------------------------------------------------------
Attachment 1: YEND_LAYOUT.png  420 kB  | Hide | Hide all
YEND_LAYOUT.png
Attachment 2: 4th_FSR1.png  60 kB  | Hide | Hide all
4th_FSR1.png
Attachment 3: 4th_FSR2.png  37 kB  | Hide | Hide all
4th_FSR2.png
ELOG V3.1.3-